OpenStudy (anonymous):

How do you solve arctan ((sqrt 3)/3)?

OpenStudy (jdoe0001):

well, you can use the Unit Circle for that, keeping in mind that \(\bf \cfrac{1}{\sqrt{3}} \textit{ simplified }\cfrac{1}{\sqrt{3}} \times \cfrac{\sqrt{3}}{\sqrt{3}} \implies \cfrac{\sqrt{3}}{\sqrt{3^2}} \implies \cfrac{\sqrt{3}}{3}\\ \textit{thus } tan^{-1}\left(\cfrac{\sqrt{3}}{3}\right) \implies tan^{-1}\left(\cfrac{1}{\sqrt{3}}\right)\)

OpenStudy (jdoe0001):

so the angle whose tangent yields that, is the angle :) so tan = sin/cos thus sine/cosine must yield \(\cfrac{1}{\sqrt{3}}\)